OpenStudy (anonymous):

Find the values of c such that the area of the region bounded by the parabolas y = 16x2 − c2 and y = c2 − 16x2 is 144

OpenStudy (psymon):

Well, these functions are completely opposite. One is the upside down version of the other. What this also means is that each function contributes half of the toal area. So if you can find a value of c such that integrating ONE of the above functions gives an area of 72, then you have your answer.

OpenStudy (anonymous):

Okay. So.. you set the two equal to each other to find the bounds, right ?

OpenStudy (psymon):

Right.

OpenStudy (anonymous):

so you would get x=plus or minus \[\sqrt{(1/16)c ^{2}}\]

OpenStudy (anonymous):

Yes ?

OpenStudy (psymon):

You put it in a different way than I would of xD \[x=\pm \frac{ c }{ 4 } \]

OpenStudy (anonymous):

Oh. Yours is prettier, lol. But are they the same thing, how does one simplify to the other ? 0.0

OpenStudy (psymon):

Well, when you did the square root you forgot the plus/minus xD But when we factored we got 2(16x^2 - c^2), which is a difference of squares that I can factor to 2(4x+c)(4x-c). From here, I just set each factor = 0 4x + c = 0 4x - c = 0

OpenStudy (anonymous):

Ooooh, okay. Gotcha, cool. So, once you get that, you would set the integration of the subtraction of the two formulas equal to 144 right ?

OpenStudy (psymon):

Well, we don't really want to do that, it'll make our work more complex. What we have our opposite function, one as a parabola that opens down and one as a parabola that opens up. Because of the symmetry, each function will contribute to half of the area, so 72. So all we really need to do is deal with one function